Consider the following product structure: Each A consists of 2 B’s and 1 C. Each B consists of 3 D’s and 2 E’s. Each C consists of 5 D’s and 3 E’s. How many D’s are necessary to produce 400 A’s?
Answer #1

400 A's are required

To produce 400 A's, 400*2 = 800 B's and 400*1=400 C's are required

To produce 800 B's, 800*3= 2400 D's and 800*2= 1600 E's are required

To produce 400 C's 400*5 = 2000 D's and 400*3 = 200 E's are required

No of D's are necessary to produce 400 A’s = 2400 + 2000 = 4400